Please show how to do these enumeration questions by showing the work.

What is the dilution factor if you add 625 µL of culture to 4.375 mL of dilution medium?

You count the 10-3 dilution of a yeast suspension using the 1/25 mm2 boxes : you find 45 cells in box 1, 15 cells in box 2, 42 cells in box 3, 23 cells in box 4, and 26 cells in box 5.

a) What is the original concentration of the yeast culture? Correct unit!

b) How many µL of which dilution do you need to spread on an agar plate to get between 100 and 200 colonies?

You perform the following dilutions: #1: 20ml stock culture into 80ml media

#2: 4ml of the #1 dilution into 196ml media

What is the dilution factor of dilution #2 in relation to the stock culture?

You count the 10-2 dilution of a yeast suspension using the 1/25 mm2 boxes : you find 40 cells in box 1, 15 cells in box 2, 46 cells in box 3, 20 cells in box 4, and 29 cells in box 5.

a) What is the original concentration of the yeast culture? Correct unit!

b) How many µL of which dilution do you need to spread on an agar plate to get between 100 and 200 colonies?

You counted 20 bacterial colonies on a plate where you plated 200 µL of the 10-4 dilution of your original yeast culture. How many CFU/mL were in the original culture?

You counted 40 bacterial colonies on a plate where you plated 400 µL of the 10-4 dilution of your original yeast culture. How many CFU/mL were in the original culture?

7) How much water and how much culture do you need to mix if you want to make a 1:6 dilution of your sample with a final volume of 800 µL?
Answer #1

What is the dilution factor if you add 625 µL of culture to 4.375 mL of dilution medium?

=

Dilution factor= (aliquot vol + diluent vol)/aliquot vol

Dilution factor= (625 + 4375)/625 = 8

You count the 10-3 dilution of a yeast suspension using the 1/25 mm2 boxes : you find 45 cells in box 1, 15 cells in box 2, 42 cells in box 3, 23 cells in box 4, and 26 cells in box 5.

a) What is the original concentration of the yeast culture? Correct unit!

cells/mL= average cell count x df x 10^4

cells/mL=[(45+15+42+23+26)/5] x 10^3 x 10^4 = 3 x 10^8

= 3 x 10^8 cells/mL

b) How many µL of which dilution do you need to spread on an agar plate to get between 100 and 200 colonies?

CFU/mL= no. of colonies/(Df x volume)

No. of colonies/(CFU/mL x DF)= Vol

Vol=200/(3 x 10^8 x 10^-3)=0.00066 mL= 0.66uL